# The Structure of Hypergraphs Arising in Cellular Mobile Communication Systems

### v3

@@ Abstract
An assumption that has often been used by researchers to model the interference in a wireless network is the unit disk graph model. While many theoretical results and performance guarantees have been obtained under this model, an open research direction is to extend these results to hypergraph interference models. Motivated by recent results that the worst-case performance of the distributed maximal scheduling algorithm is characterized by the interference degree of the hypergraph, in the present work we investigate properties of the interference degree of the hypergraph and the structure of hypergraphs arising from physical constraints. We show that the problem of computing the interference degree of a hypergraph is NP-hard and we prove some properties and results concerning this hypergraph invariant. We investigate which hypergraphs are realizable, i.e. which hypergraphs arise in practice, based on physical constraints, as the interference model of a wireless network. In particular, a question that arises naturally is: what is the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable? We determine this quantity for various integral and nonintegral values of the path loss exponent of signal propagation. We also investigate hypergraphs generated by line networks.
26 December 2023
Index terms — Cellular systems; hypergraph models; unit disk graphs; wireless networks; distributed algorithms; maximal scheduling; interference degree; applications of hypergraph theory.
@@ 1 Introduction
Combinatorial interference models for wireless communication networks such as conflict graphs and conflict hypergraphs have been widely studied in the literature, and the performance of various protocols and distributed algorithms for resource allocation and scheduling have been investigated using these models. The combinatorial objects that arise in this context due to geometric, graph-theoretic, or physical (SINR) constraints have additional structure that can be exploited to design efficient algorithms that are guaranteed to be close to optimal for various problems. For example, unit disk graphs have additional structure - the interference degree of a unit disk graph (defined below) is at most $5$ and this implies that the performance of the distributed maximal scheduling algorithm is a factor of at most $5$ away from optimal. An open problem in the literature that we consider in the present work is to extend the results in the literature on unit disk graphs to hypergraph models.
@@
The rest of this paper is organized as follows. Section 2 describes the system model; formal definitions of the unit disk graph model and hypergraph model are given, and the interference degree of a hypergraph is defined. Section 3 discusses related work. In Section 4, we show that the problem of computing the interference degree of a hypergraph is NP-hard and determine the interference degree of some structured hypergraphs. In Section 5 we prove some preliminary results, which are used in Section 6, Section 7, Section 8, and Section 9 to prove that certain hypergraphs are not realizable when $\alpha=4,3,2$ and $1$, respectively. In Section 10, these results are extended to nonintegral values of the path loss exponent $\alpha$. In Section 11 some results concerning the hypergraphs of line networks are obtained. Section 12 contains concluding remarks.
@@ 2.2 Signal Propagation Model and Path Loss Exponents
The path loss exponent plays a pivotal role in characterizing the attenuation of a wireless signal as it traverses through the radio channel. This exponent, denoted here by $\alpha$, quantifies the rate at which the signal strength diminishes with increasing distance from the transmitter. Understanding the path loss exponent is crucial in designing and optimizing wireless communication systems, as it directly influences signal coverage, range, and overall system performance. The path loss exponent is particularly significant in the deployment of wireless networks, where variations in the propagation environment necessitate a nuanced consideration of signal attenuation for efficient communication and reliable connectivity.
While signal propagation effects exhibit complexity in real channels and vary across different environments, a simplified path loss model is often employed for the analysis of tradeoffs and system design. This model expresses the received signal strength as a function of distance. In the scenario where $s_{i}$ acts as the transmitter, $s_{j}$ serves as the receiver, and $d(s_{i},s_{j})$ represents the distance between them, the received signal strength at $s_{j}$ is defined as $\frac{1}{d(s_{i},s_{j})^{\alpha}}$, with $\alpha$ denoting the path loss exponent. This model has been used in the literature on hypergraph interference models (defined below), where $\alpha$ was taken to equal $4$ in [39] [43] [42]. Typical values of the path loss exponent for various environments are presented in Table 1; cf [11] [23].
Environment $\alpha$ range
@@
Table 1: Typical path loss exponents. Table courtesy of [23, Section 2.5].
@@ 2.3 The Hypergraph of a Wireless Network
Consider a wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$, where $S=\{s_{1},s_{2},\ldots,s_{n}\}$ is a set of stations (or transceivers), $\alpha$ is the path-loss exponent, and $\beta$ is the reception threshold. The parameter $\beta$ is sometimes also called the SINR threshold for successful reception [43] (though we ignore noise in the present model), and refers to the maximum interference that is tolerable. We shall identify the symbol $s_{i}$, used for a station, with its location in the Euclidean plane, so that $S\subseteq\mathbb{R}^{2}$. The value of the path loss exponent $\alpha$ typically varies between $1.6$ and $6.5$ [23].
@@
Given a wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$, a subset $W\subseteq S$ is called a forbidden set for the wireless network $\mathcal{A}$ if the interference is intolerable at some station in $W$ when all the other stations in $W$ are active. In other words, a subset $W$ of stations is forbidden if there exists a station $w\in W$ such that
$E(W-\{w\},w)=\sum_{s\in W-\{w\}}\frac{1}{d(s,w)^{\alpha}}\geq\beta.$
@@
Let $\mathcal{A}=\langle S,\alpha,\beta\rangle$ be a wireless network, where $S$ is a set of points in $\mathbb{R}^{2}$, $\alpha$ is the path loss exponent, and $\beta$ is the reception threshold. The hypergraph generated by $\mathcal{A}$ is defined to be the hypergraph $H=(S,\mathcal{E})$, where $\mathcal{E}$ is the collection of all minimal forbidden sets.
@@
2.4 The Interference Degree of a Hypergraph
In this subsection, we recall from the literature [18] the definition of the interference degree of a hypergraph, which characterizes the worst-case performance of the distributed maximal scheduling algorithm.
@@ 2.4 The Interference Degree of a Hypergraph
Let $\mathcal{I}(H)$ denote the set of all independent sets of the hypergraph $H$.
@@ 2.5 Notation and Terminology
If $S$ is a finite set, then $S^{(r)}$ denotes the collection of all $r$-subsets of $S$. A hypergraph $(S,\mathcal{E})$ is said to be $r$-uniform if the size of each hyperedge is $r$, i.e. if $\mathcal{E}\subseteq S^{(r)}$. Thus, a $2$-uniform hypergraph is just a graph. The complete $r$-uniform hypergraph, denoted by $K_{n}^{(r)}$, is defined to be the hypergraph with vertex set $[n]$ and with the edge set equal to the collection $[n]^{(r)}$ of all $r$-subsets of $[n]$.
@@ 3 Related Work
For example, the unit disk graph assumption has been well-studied by researchers in the sensor networks community. In [37], an algorithm is given that has a performance guarantee of $3$ for the maximum independent set problem for unit disk graphs, even though this problem is hard to approximate for general graphs. In [37], an online algorithm for vertex coloring is given that has a competitive ratio of $6$ for unit disk graphs. The interference degree of a unit disk graph is at most $5$, and so the maximal scheduling algorithm is a factor of at most $5$ away from optimal for unit disk graphs [13] [15]. A distributed algorithm for admission control known as the scaled clique constraints are a factor of at most $2.2$ away from optimal for unit disk graphs [25].
@@ 4 The Interference Degree of Structured Hypergraphs
Given that the worst-case performance of the maximal scheduling algorithm under the hypergraph interference model is characterized by the interference degree of the hypergraph, in the present section we investigate certain properties of this hypergraph invariant. We show that the problem of computing the interference degree of a hypergraph is NP-hard, and we prove some basic properties of this hypergraph invariant and compute the interference degree of certain structured hypergraphs. These results extend or generalize the previous results in the literature on graph models to hypergraph models.
@@
Proof: See [21].
@@
Proof: See [21].
@@
Proof: See [21].
@@
Proof: See [21].
@@
Proof: See [21].

### v4

@@ Abstract
An assumption that researchers have often used to model interference in a wireless network is the unit disk graph model. While many theoretical results and performance guarantees have been obtained under this model, an open research direction is to extend these results to hypergraph interference models. Motivated by recent results that the worst-case performance of the distributed maximal scheduling algorithm is characterized by the interference degree of the hypergraph, in the present work we investigate properties of the interference degree of the hypergraph and the structure of hypergraphs arising from physical constraints. We show that the problem of computing the interference degree of a hypergraph is NP-hard and we prove some properties and results concerning this hypergraph invariant. We investigate which hypergraphs are realizable, i.e. which hypergraphs arise in practice, based on physical constraints, as the interference model of a wireless network. In particular, a question that arises naturally is: what is the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable? We determine this quantity for various integral and nonintegral values of the path loss exponent of signal propagation. We also investigate hypergraphs generated by line networks.
Index terms — Cellular systems; hypergraph models; unit disk graphs; wireless networks; applied algorithms; distributed algorithms; maximal scheduling; interference degree; applications of hypergraph theory.
@@ 1 Introduction
Combinatorial interference models for wireless communication networks such as conflict graphs and conflict hypergraphs have been widely studied in the literature, and the performance of various protocols and distributed algorithms for resource allocation and scheduling have been investigated using these models. The combinatorial objects that arise in this context due to geometric, graph-theoretic, or physical constraints have additional structure that can be exploited to design efficient algorithms with near-optimal performance for various problems. Physical constraints can include the requirement that the SINR (Signal-to-Interference-plus-Noise Ratio) exceed a certain threshold. For example, unit disk graphs have additional structure - the interference degree of a unit disk graph (defined below) is at most $5$ and this implies that the performance of the distributed maximal scheduling algorithm is a factor of at most $5$ away from optimal. An open problem in the literature that we consider in the present work is to extend the results in the literature on unit disk graphs to hypergraph models.
@@
The rest of this paper is organized as follows. Section 2 describes the system model; formal definitions of the unit disk graph model and hypergraph model are given, a distributed maximal scheduling algorithm from the literature is recalled, and it is explained that its worst-case performance is characterized by the interference degree of the hypergraph. Section 3 discusses related work. In Section 4, we show that the problem of computing the interference degree of a hypergraph is NP-hard and determine the interference degree of some structured hypergraphs. In Section 5 we prove some preliminary results, which are used in Section 6, Section 7, Section 8, and Section 9 to prove that certain hypergraphs are realizable or nonrealizable when the path loss exponent $\gamma=4,3,2$ and $1$, respectively. In Section 10, these results are extended to nonintegral values of the path loss exponent $\gamma$. In Section 11 some results concerning the hypergraphs of line networks are obtained. Section 12 contains concluding remarks. To facilitate understanding, we provide in Table 1 a list of all symbols used along with their corresponding meanings and the section where they are first defined.
Symbol Meaning Section
$S$ A set of $\{s_{1},\ldots,s_{n}\}\subseteq\mathbb{R}^{2}$ of stations in the Euclidean plane 2.1
$G=(V,E)$ a simple, undirected graph 2.1
$\alpha(G)$ independence number of the graph $G$ 2.1
$\Gamma(v)$ the set of neighbors of vertex $v$ 2.1
$G[W]$ subgraph of $G$ induced by $W$ 2.1
$\sigma(G)$ interference degree of the graph $G$ 2.1
$K_{1,r}$ the star graph, or the $2$-uniform star hypergraph 2.1
$\gamma$ path loss exponent 2.2
$d(a,b)$ Euclidean distance between points $a$ and $b$ 2.2
$H=(V,\mathcal{E})$ hypergraph with vertex set $V$ and edge set $\mathcal{E}$ 2.3
$\beta$ SINR threshold for successful reception 2.3
$\mathcal{A}=\langle S,\gamma,\beta\rangle$ wireless network, with set of stations $S$, path loss exponent $\gamma$, and reception threshold $\beta$ 2.3
$E(W,w)$ energy or interference at $w$ due to transmission by stations in $W$ 2.3
$\tau=(\tau(v):v\in V)$ QoS requirement or demand vector 2.4
$\mathcal{I}(H)$ collection of all independent sets of the hypergraph $H$ 2.4
$A=[a_{ij}]$ vertex-independent set incidence matrix of a hypergraph 2.4
$\chi_{f}(H,\tau)$ fractional chromatic number of the weighted hypergraph $(H,\tau)$ 2.4
$N(i)$ set of neighbors in a hypergraph of vertex $i$ 2.4
$\Delta_{ij},\Delta_{i}^{\prime},\Delta_{i}^{\prime\prime}$ intermediate variables for interference degree calculation 2.4
$B(H,\tau)$ an upper bound on $\chi_{f}(H,\tau)$ 2.4
$\eta(H)$ worst-case performance of the maximal scheduling algorithm 2.4
$\sigma(H)$ interference degree of the hypergraph $H$ 2.4
$[N]$ the set $\{1,2,\ldots,N\}$ 2.5
$S^{(r)}$ the collection of all $r$-subsets of $S$ 2.5
$K_{n}^{(r)}$ the complete $r$-uniform hypergraph 2.5
$\theta_{i}$ angle between stations on a unit circle 6
$\delta_{i}$ translated version of angle between stations on a unit circle 6
$f(\delta_{1})$ cumulative interference at a station due to transmissions by its two closest neighbors 6
$g(\delta)$ a translation of the function $f(\delta_{1})$ 6
$r(\gamma)$ largest star hypergraph that is realizable for path loss exponent $\gamma$ 10
$\mathcal{U}_{k}$ a set of $k$ stations placed uniformly on the unit circle 10
$u(\gamma)$ largest $k$ such that $\mathcal{U}_{k}$ is feasible 10
$P_{n}$ the path graph on $n$ vertices, or the $2$-uniform path hypergraph on $n$ vertices 11
Table 1: Notation index in the order of first occurrence in paper
@@ 2.2 Signal Propagation Model and Path Loss Exponents
The path loss exponent plays a pivotal role in characterizing the attenuation of a wireless signal as it traverses through the radio channel. This exponent, denoted here by $\gamma$, quantifies the rate at which the signal strength diminishes with increasing distance from the transmitter. Understanding the path loss exponent is crucial in designing and optimizing wireless communication systems, as it directly influences signal coverage, range, and overall system performance. The path loss exponent is particularly significant in the deployment of wireless networks, where variations in the propagation environment necessitate a nuanced consideration of signal attenuation for efficient communication and reliable connectivity.
While signal propagation effects exhibit complexity in real channels and vary across different environments, a simplified path loss model is often employed for the analysis of tradeoffs and system design. This model expresses the received signal strength as a function of distance. It is assumed that the power level is the same for all transmitters and that the signal strength degrades uniformly in all directions. In the scenario where $s_{i}$ acts as the transmitter, $s_{j}$ serves as the receiver, and $d(s_{i},s_{j})$ represents the distance between them, the received signal strength at $s_{j}$ is defined as $\frac{1}{d(s_{i},s_{j})^{\gamma}}$, with $\gamma$ denoting the path loss exponent. This model has been used in the literature on hypergraph interference models (defined below), where $\gamma$ was taken to equal $4$ in [41] [44] [45]. Typical values of the path loss exponent for various environments are presented in Table 2; cf [11] [25]. In most real environments, the path loss exponent $\gamma>1$, although $\gamma<1$ in certain indoor settings. In this paper, it is assumed throughout that $\gamma>0$.
Environment $\gamma$ range
@@
Table 2: Typical path loss exponents. Table courtesy of [25, Section 2.5].
@@ 2.3 The Hypergraph of a Wireless Network
Consider a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$, where $S=\{s_{1},s_{2},\ldots,s_{n}\}$ is a set of stations (or transceivers), $\gamma$ is the path-loss exponent, and $\beta$ is the reception threshold. The parameter $\beta$ is sometimes also called the SINR threshold for successful reception [44] (though we ignore noise in the present model), and refers to the maximum interference that is tolerable. We shall identify the symbol $s_{i}$, used for a station, with its location in the Euclidean plane, so that $S\subseteq\mathbb{R}^{2}$. The value of the path loss exponent $\gamma$ typically varies between $1.6$ and $6.5$ [25].
@@
Given a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$, a subset $W\subseteq S$ is called a forbidden set for the wireless network $\mathcal{A}$ if the interference is intolerable at some station in $W$ when all the other stations in $W$ are active. In other words, a subset $W$ of stations is forbidden if there exists a station $w\in W$ such that
$E(W-\{w\},w)=\sum_{s\in W-\{w\}}\frac{1}{d(s,w)^{\gamma}}\geq\beta.$
@@
Let $\mathcal{A}=\langle S,\gamma,\beta\rangle$ be a wireless network, where $S$ is a set of points in $\mathbb{R}^{2}$, $\gamma$ is the path loss exponent, and $\beta$ is the reception threshold. The hypergraph generated by $\mathcal{A}$ is defined to be the hypergraph $H=(S,\mathcal{E})$, where $\mathcal{E}$ is the collection of all minimal forbidden sets.
@@
2.4 A Distributed Maximal Scheduling Algorithm and The Interference Degree of a Hypergraph
In this subsection, we recall from the literature a certain distributed maximal scheduling algorithm [38] and explain how its worst-case performance is characterized by the interference degree of the hypergraph [18].
Consider a network whose entities compete for some resource, and the competition is modeled by a conflict hypergraph $H=(V,\mathcal{E})$, where the vertex set $V$ represents the set of entities, and each hyperedge $E\in\mathcal{E}$ represents a minimal collection of entities that cannot all be assigned the resource at the same time due to said conflict or competition. Furthermore, each entity $v\in V$ specifies a fraction $\tau(v)$ of each unit of time that it would like to use the resource. The objective is to determine whether the demand vector $\tau$ can be satisfied. For example, in a wireless network, the entities are the wireless links (or stations), and links in the same vicinity contend for the shared wireless medium due to interference. The quality-of-service (QoS) requirement is specified in terms of a minimum bandwidth requirement for each link, which translates into a fraction of each unit of time when the link should be active. Each hyperedge represents a minimal forbidden set of links. The admission control problem is to determine if the demand vector is feasible, i.e. whether all the demands can be satisfied. It is preferable that this decision be made in a distributed manner, i.e. using only localized information.
The admission control problem can be reformulated as follows. Let $H=(V,\mathcal{E})$ be a conflict hypergraph, with $V=\{v_{1},\ldots,v_{n}\}$, and let $\tau=(\tau(v):v\in V)$ be a demand vector. Let $\mathcal{I}(H)=\{I_{1},\ldots,I_{L}\}$ be the collection of all independent sets of $H$. Let $A=[a_{ij}]$ be the $0,1$-valued $N\times L$ vertex-independent set incidence matrix of $H$; thus $a_{ij}=1$ if $v_{i}\in I_{j}$, and $a_{ij}=0$ otherwise. The fractional chromatic number of the weighted hypergraph $(H,\tau)$, denoted by $\chi_{f}(H,\tau)$, is the value of the linear program: min $1^{T}t$ subject to $Mt\geq\tau$, $t\geq 0$. It can be seen that a demand vector $\tau$ is feasible if and only if $\chi_{f}(H,\tau)\leq 1$. Next, we shall give a sufficient condition for admission control that can be implemented in a distributed manner.
@@ 2.4 A Distributed Maximal Scheduling Algorithm and The Interference Degree of a Hypergraph
Theorem 2.
[38] Let $H$ be a hypergraph, with $\Delta_{ij}=\Delta_{ij}(H)$ as defined above. A sufficient condition for a demand vector $\tau$ to be feasible is that
$\tau(v_{i})+\sum_{j\neq i}\left\{\Delta_{ij}\tau(v_{j})\right\}\leq 1,\forall i\in[N].$
Theorem 2 gives a distributed mechanism for admission control and yields a so-called maximal scheduling algorithm. The actual resource requirement for satisfying demand $\tau$ is given by $\chi_{f}(H,\tau)$. An upper bound estimate for this resource requirement, as given by Theorem 2, is
$B(H,\tau)=\max_{i\in[N]}\left\{\tau(v_{i})+\sum_{j\neq i}\left\{\Delta_{ij}\tau(v_{j})\right\}\right\}.$
We know $\chi_{f}(H,\tau)\leq B(H,\tau)$. The distributed maximal scheduling algorithm is suboptimal because it can overestimate the resource requirements needed to satisfy a given demand vector in the sense that it is possible that $\chi_{f}(H,\tau)\leq 1$, but $B(H,\tau)>1$. In the worst case, this sufficient condition for feasibility overestimates the resource requirements by a factor equal to $\eta(H)$, where
$\eta(H)=\sup_{\tau\neq 0}\frac{B(H,\tau)}{\chi_{f}(H,\tau)}.$
It turns out this worst case performance is exactly equal to the interference degree of the hypergraph, defined next.
@@
The following theorem highlights the importance of the interference degree $\sigma(H)$:
Theorem 4.
[18] Let $H=(V,\mathcal{E})$ be a conflict hypergraph. The worst-case performance $\eta(H)$ of the distributed maximal scheduling algorithm is equal to the interference degree $\sigma(H)$.
@@ 2.5 Notation and Terminology
If $S$ is a finite set, then $S^{(r)}$ denotes the collection of all $r$-subsets of $S$. A hypergraph $(S,\mathcal{E})$ is said to be $r$-uniform if the size of each hyperedge is $r$, i.e. if $\mathcal{E}\subseteq S^{(r)}$. Thus, a $2$-uniform hypergraph is just a graph. The complete $r$-uniform hypergraph of order $n$, denoted by $K_{n}^{(r)}$, is defined to be the hypergraph with vertex set $[n]$ and with the edge set equal to the collection $[n]^{(r)}$ of all $r$-subsets of $[n]$.
@@ 3 Related Work
For example, the unit disk graph assumption has been well-studied by researchers in the sensor networks community. In [39], an algorithm is given that has a performance guarantee of $3$ for the maximum independent set problem for unit disk graphs, even though this problem is hard to approximate for general graphs. In [39], an online algorithm for vertex coloring is given that has a competitive ratio of $6$ for unit disk graphs. The interference degree of a unit disk graph is at most $5$, and so the maximal scheduling algorithm is a factor of at most $5$ away from optimal for unit disk graphs [13] [15]. A distributed algorithm for admission control known as the scaled clique constraints is a factor of at most $2.2$ away from optimal for unit disk graphs [27].
@@ 4 The Interference Degree of Structured Hypergraphs
In the present section we investigate certain properties of a hypergraph invariant - the interference degree. We show that the problem of computing the interference degree of a hypergraph is NP-hard, and we prove some basic properties of this hypergraph invariant and compute the interference degree of certain structured hypergraphs. These results extend or generalize the previous results in the literature on graph models to hypergraph models.
@@
Proof: The proof is by reduction from the independence number problem. The independence number $\alpha(G)$ of a graph $G$ is the maximum number of vertices that are pairwise nonadjacent. The problem of computing the independence number of a graph is NP-hard [23]; in fact, even approximating the independence number to within a constant factor is computationally hard. Given a graph $G=(V,E)$, construct an augmented graph $G^{\prime}$, which is obtained from $G$ by joining a new vertex $x$ to each vertex of $G$. Thus, $G^{\prime}$ has $|V|+1$ vertices and $|E|+|V|$ edges. The interference degree $\sigma(G^{\prime})$ of the augmented graph $G^{\prime}$ is equal to the independence number $\alpha(G)$. Hence, $\alpha(G)\geq k$ if and only if $\sigma(G^{\prime})\geq k$.
@@
Proof: Given any graph $G=(V,E)$, construct a $2$-uniform hypergraph $H=(V,E)$ with the same set of edges as $G$. Then, the interference degree of the hypergraph $H$ is equal to the interference degree of the graph $G$ [18, p. 2955]. Hence, the problem of computing the interference degree of a graph is a special case of the problem of computing the interference degree of a hypergraph. The assertion follows from Theorem 5.
@@
Proof: Consider the complete $r$-uniform hypergraph $K_{n}^{(r)}$, which has vertex set $[n]$ and edge set $[n]^{(r)}$. By Definition 3, we need to determine $\Delta_{i}^{\prime}$ and $\Delta_{i}^{\prime\prime}$ ($i\in[n]$). By symmetry, these values are independent of $i$. So fix $i\in[n]$. Since the hypergraph is complete, each vertex is in the neighborhood set of all other vertices, i.e. $N(i)=V-\{i\}$. Each hyperedge has size $r$, whence $\Delta_{ij}=\frac{1}{r-1}$, for all $j\neq i$. The hyperedges are exactly the $r$-subsets of the vertex set, which implies that the maximum size of an independent set $J$ in $N(i)$ is $r-1$. Hence,
$\Delta_{i}^{\prime}=\max_{J\subseteq N(i):J\in\mathcal{I}(H)}\sum_{j\in J}\Delta_{ij}=(r-1)\frac{1}{r-1}=1.$
The maximum size of an independent set $J$ in $N(i)$ such that $J\cup\{i\}$ is an independent set is $r-2$. Hence,
$\Delta_{i}^{\prime\prime}=\max_{J\subseteq N(i):J\cup\{i\}\in\mathcal{I}(H)}1+\sum_{j\in J}\Delta_{ij}=1+\frac{r-2}{r-1}=2-\frac{1}{r-1}.$
By Definition 3,
$\sigma(H)=\max\{1,2-\frac{1}{r-1}\}=2-\frac{1}{r-1}.$
@@
Proof: Let $H=(V,\mathcal{E})$ be a hypergraph. By Definition 3, $\Delta_{ij}\leq 1$, for all $i,j\in V$. Hence, for $i\in V$,
$|\Delta_{i}^{\prime}|\leq\max_{J\subseteq N(i):J\in\mathcal{I}(H)}|J|\leq\alpha(H),$
and
$|\Delta_{i}^{\prime\prime}|\leq\max_{J\subseteq N(i):J\cup\{i\}\in\mathcal{I}(H)}1+|J|\leq\alpha(H).$
Hence, $\sigma(H)\leq\alpha(H)$.
@@
Proof: By definition of $\Delta_{i}^{\prime\prime}$, we have $\sigma(H)\geq 1$. If the edge set $\mathcal{E}$ is empty, then the neighborhood sets $N(i)$ are empty, and so $\Delta_{i}^{\prime\prime}=1$ for all $i$ and $\sigma(H)=1$. If the edge set is nonempty, then the maximum size of an independent set in $H$ is at most $n-1$. By Lemma 8, $\sigma(H)\leq n-1$. The star hypergraph $K_{1,n-1}$ attains the upper bound, and the complete $2$-uniform hypergraph $K_{n}^{(2)}$ attains the lower bound.